The transverse diaplacement for a wave on a string traveling in the +x direction is y(x,t) = (0.0090) sin (66.8m^-1)x-(310s^-1)t) what is it’s a) velocity b)wavelength c)frequency d)period e)amplitude (in units)
Answer #1

compare given equation with

Y=Asin(kx-wt)

amplitude

A=0.009 m

wave number

K=66.8 m-1

angular frequency

W=310 rad/sec

wave length

lambda =2pi/K =2pi/66.8=0.094 m

frequency

f=W/2pi =310/2pi =49.36 Hz

a)

velocity

V=f*lambda =49.36*0.094

V=4.64 m/s

b)

wavelength

lambda =0.094 m

c)

frequency

f=49.36 Hz

d)

period

T=1/f =1/49.36=0.02026 s

e)

amplitude

A=0.009 m